HUBBLE SPACE TELESCOPE DAILY REPORT #1198 PERIOD COVERED: 0000Z Daily Status Report as of 221/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5629 (WF/PC-2 Cycle 4 Photometric Calibration Monitor 1C: 4-HIP STD in F170W) The WF/PC-2 was used to make photometric calibration observations of the ultraviolet standard star GRW+70D582. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5419 (A Search for Very Low-Luminosity Active Galactic Nuclei in Nearby Galaxies) The WF/PC-2 was used to make observations of galaxy NGC-4036, which is known to have a ultraviolet bright, low-luminosity, active galactic nuclei. The observations were completed as planned, and no problems were noted. 1.3 Completed WF/PC-2 5659 (WF/PC-2 Cycle 4 Calibration: CTE Dither Test: Part 2) The WF/PC-2 was used to make CTE (Charges Transfer Efficiency) Dither Calibration images by making observations of the large globular cluster Omega-Cen. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed FOS/BL 5658 (White Dwarf Calibration) The Faint Object Spectrograph (f/96) was used to make absolute sensitivity calibrations of the detectors using standard white dwarf star HZ43. The observations were executed as planned, and no anomalies were noted. 1.5 Completed WF/PC-2 5446 (The Nuclei of Nearby S0 and Spiral Galaxies) The WF/PC-2 was used to make observations of the core region of galaxy NGC-4634 to provide statistics on the nuclear cups, stellar condensations, and inner stellar disks and dust lanes. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 5667 (Binaries in Globular Clusters NGC-6441 Image (FOS 37) Carry-over Observation) The WF/PC-2 was used to make observations of the globular cluster NGC-6441 using several different filters. These will be "early-acq" images that will be used to locate objects with a ultraviolet excess for later observations. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed FOC/96 5126 (Mass Exchange Binaries (FOS 34) V1343-AQL: Cycle 4) The Faint Object Camera (f/96) was used used to make observations of the mass exchange binary star V1343-Aql. These observations were taken in parallel with images of the jets in SS433 (aka, V1343 Aql) to look for optical and ultraviolet emissions. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 5397 (Determination of the Extragalactic distance Scale: Cycle 4 High) The WF/PC-2 was used to make images of galaxy NGC-925 to be used in measuring the periods of Cepheid variable stars within this galaxy. The images were received and displayed, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 7 FGS 1 and 2 guide star acquisition failed to course track at 220/2025Z during ZOE (HSTAR 4866). The roll maneuver update may not have corrected all of the errors accumulated during the T2 slew. Scheduled Reacquisitions : 7 Successful : 5 Two failed to course track (HSTAR 4866) Related to HSTAR 4866. Subsequent acquisition was successful. 2.2 FHST Updates: Scheduled :16 Successful :16 2.3 Significant HSTARs: None 2.4 Operations Notes: DOB Move requested by the FOC science team Last week a one-step move of the COSTAR FOC M1 mirror was performed. The objective, as stated by the COSTAR FOC team, was to remove a small amount of residual coma in the FOC point spread function by executing a small (one-step) x-tilt change of the mirror before selected FOC observations which began on day 216. This move changed the x-tilt from 61.9 arc-seconds to 54.8 arc-seconds. This week, as part of the same analysis, the FOC team has requested that the DOB be moved 12 steps (about 0.2 mm) in the negative direction to adjust the FOC focus. This magnitude change is expected to have little or no effect on the other instruments using the COSTAR. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-